### v1

# On the controller form for linear hyperbolic MIMO systems with dynamic boundary conditions

Abstract
This contribution develops an algebraic approach to obtain a controller form for a class of linear hyperbolic MIMO systems, bidirectionally coupled with a linear ODE system at the unactuated boundary. After a short summary of established controller forms for SISO and MIMO ODE as well as SISO hyperbolic PDE systems, it is shown that the direct approach to state a controller form fails already for a very simple MIMO example. Next, a generalised hyperbolic controller form with different variants is proposed and a new flatnesss-based scheme to compute said form is presented. Therein, the system is treated in an algebraic setting where generalised polynomials with real exponents are used to describe the predictions and delays in the system. The proposed algorithm is then applied to the motivating example.
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
I Introduction
From the onset of state space based approaches, controller (canonical) forms have been a vital asset for systematic controller designs. As the controller forms for linear SISO and MIMO ODE systems, cf. [1, 2] as well for SISO hyperbolic PDE systems [3] are long established and understood, only comparatively recent works [4, 5, 6] discuss hyperbolic systems with boundaries governed by ODEs for the SISO case which this contribution extends to the MIMO case.
Following some preliminaries, where we briefly restate established controller forms and demonstrate the problems that arise if a hyperbolic controller form is to be introduced even for the most simple MIMO system, a MIMO hyperbolic controller form (HCF) is given in Section III. The main part of the contribution in Section IV then discusses how a HCF can be obtained for a specific class of systems whereupon Section V tracks the procedure for an example system.
II Preliminaries
II-A Notation
The variables $z$ and $t$ denote space and time. The first and second derivative of $x(t)$ is denoted by $\dot{x}(t)$ and $\ddot{x}(t)$, while the $n$-th derivative is given by $x^{(n)}(t)$ and the partial derivative of $x(z,t)$ w.r.t. $z$. by $\partial_{z}x(z,t)$.
To treat delay differential equations (DDEs) in an algebraic setting, we identify the operators therein with their Fourier-Laplace transforms and treat the application of the transformed operators as convolutions. Thus, we write $\mathcal{L}\{\partial_{t}\}=s$ for differentials, while the shifts $(\sigma^{\tau}x)(t)=x(t+\tau)$ with $\tau\in\mathbb{R}$ are denoted as $\mathcal{L}\{\sigma^{\tau}\}(s)=\exp(s\tau)$.
The shorthand $L^{2}([a,b])$ is used for the space of square integrable functions on $[a,b]$, while $H^{n}([a,b])$ stands for the space of functions whose $n$-th derivative belongs to $L^{2}([a,b])$ and $\mathcal{E}^{\prime}$ for Schwartz-Distributions with compact support. With $\operatorname{V}_{f}([a,b])$ denoting the total variation of a function $f$ on the interval $[a,b]$, $BV([a,b])$ contains functions of bounded variation while members of its subset $BV^{\star}([a,b])$ satisfy $\lim_{\epsilon\to 0}\operatorname{V}_{f}([b,b+\epsilon))=0$.
Furthermore, $\mathbb{R}[\partial_{t}]$ is the space of polynomials in $\partial_{t}$ with coefficients in $\mathbb{R}$ and $\mathbb{R}(\partial_{t})$ consists of rational expressions in $\partial_{t}$. Regarding a polynomial matrix $X(\partial_{t})\in\mathbb{R}^{n\times n}[\partial_{t}]$, we denote the leading column coefficient matrix (LCCM) w.r.t. $\partial_{t}$ by $\operatorname{lccm}_{\partial_{t}}{\mathbf{X}(\partial_{t})}$ and say that $\mathbf{X}(\partial_{t})$ is column reduced w.r.t. $\partial_{t}$ if $\operatorname{lccm}_{\partial_{t}}{\mathbf{X}(\partial_{t})}$ has full generic rank. Finally, we call $\mathbf{X}(\partial_{t})$ unimodular if its determinant is independent of $\partial_{t}$.
II-B Established Controller Forms
II-B1 ODE SISO systems
For this class, [1] defines the controller canonical form
$\dot{\eta}^{\mathrm{O}}_{1}(t)$ $=\eta^{\mathrm{O}}_{2}(t)$ (1a)
$\;\;\vdots$
$\dot{\eta}^{\mathrm{O}}_{n-1}(t)$ $=\eta^{\mathrm{O}}_{n}(t)$ (1b)
$\dot{\eta}^{\mathrm{O}}_{n}(t)$ $=-\bm{a}^{\intercal}\bm{\eta}^{\mathrm{O}}(t)+u(t)$ (1c)
with state $\bm{\eta}^{\mathrm{O}}(t)=(\eta^{\mathrm{O}}_{1}(t),\dotsc,\eta^{\mathrm{O}}_{n}(t))^{\intercal}\in\mathbb{R}^{n}$, coefficients $\bm{a}\in\mathbb{R}^{n}$ and input $u(t)\in\mathbb{R}$ that exists for every controllable system. Note that the “output” $\eta^{\mathrm{O}}_{1}(t)$ of the integrator chain is a flat output of the system.
II-B2 ODE MIMO systems
Extending the SISO framework to systems with inputs $\bm{u}(t)\in\mathbb{R}^{m}$, [2] states a controller form [^1] As these forms are not unique in the MIMO case, the ancillary adjective canonical will be omitted from now on. Moreover, note that (2c) explicitly allows for all inputs to occur in the “input” of the subsystems integrator chain with linear factors $\neq 1$ as long as $\mathbf{B}=(\bm{b}_{1},\dotsc,\bm{b}_{m})^{\intercal}$ has full rank. , with subsystems
$\dot{\eta}^{\mathrm{O}}_{i,1}(t)$ $=\eta^{\mathrm{O}}_{i,2}(t)$ (2a)
$\;\;\vdots$
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}-1}(t)$ $=\eta^{\mathrm{O}}_{i,\nu_{i}}(t)$ (2b)
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=-\sum_{k=1}^{m}\bm{a}_{k}^{\intercal}\bm{\eta}^{\mathrm{O}}_{k}(t)+\bm{b}_{i}^{\intercal}\bm{u}(t)$ (2c)
for $i\in\{1,\dotsc,{m}\}$, with the Kronecker index $\nu_{i}$ and the state vector $\bm{\eta}^{\mathrm{O}}_{i}(t)=(\eta^{\mathrm{O}}_{i,1}(t),\dotsc,\eta^{\mathrm{O}}_{i,\nu_{i}}(t))^{\intercal}\in\mathbb{R}^{\nu_{i}}$ of the $i$-th subsystem. Similarly to the SISO case, the MIMO controller form state $\bm{\eta}^{\mathrm{O}}(t)\in\mathbb{R}^{\nu_{\Sigma}}$ with $\nu_{\Sigma}\coloneq\sum_{i=1}^{m}\nu_{i}$ and input $\bm{u}(t)$ can be easily parametrised by the tuple $(\eta^{\mathrm{O}}_{1,1}(t),\dotsc,\eta^{\mathrm{O}}_{m,1}(t))$.
II-B3 hyperbolic SISO systems
Initially, the work [3] introduced a controller form for boundary controlled hyperbolic systems that takes the form of decoupled transport systems. These results were then extended to dynamic boundary conditions in [4, 5], yielding the controller form
$\dot{\eta}^{\mathrm{O}}_{1}(t)$ $=\eta^{\mathrm{O}}_{2}(t)$ (3a)
$\;\;\vdots$
$\dot{\eta}^{\mathrm{O}}_{n-1}(t)$ $=\eta^{\mathrm{O}}_{n}(t)$ (3b)
$\dot{\eta}^{\mathrm{O}}_{n}(t)$ $=\eta^{\mathrm{P}}(\check{\tau},t)$ (3c)
[…]
$\eta^{\mathrm{P}}(\hat{\tau},t)$ $=-\operatorname{\mathcal{S}}_{a^{\star}}\eta^{\mathrm{P}}(\cdot,t)-\bm{a}^{\intercal}\eta^{\mathrm{O}}(t)+u(t).$ (3e)
[…]
II-C Benchmark System
Consider a system of two vibrating strings
$\partial_{t}^{2}x_{1}(z,t)-\partial_{z}^{2}x_{1}(z,t)$ $=0\quad z\in(-\alpha,0)$ (4a)
$\partial_{t}^{2}x_{2}(z,t)-\partial_{z}^{2}x_{2}(z,t)$ $=0\quad z\in(0,\beta),$ (4b)
coupled with a point mass at $z=0$ and thus subject to the dynamic boundary conditions
$x_{1}(0,t)=x_{2}(0,t)$ $=\xi(t)$ (5a)
$m\ddot{\xi}(t)$ $=-\partial_{z}x_{1}(0,t)+\partial_{z}x_{2}(0,t),$ (5b)
where $\xi(t)$ denotes the mass position as well as
$\partial_{z}x_{1}(-\alpha,t)$ $=-u_{1}(t)$ (5c)
$\partial_{z}x_{2}(\beta,t)$ $=u_{2}(t),$ (5d)
with state [^2] In detail, $\bm{x}(\cdot,t)\!=\!(\xi(t),\dot{\xi}(t),x_{1}(\cdot,t),\partial_{t}x_{1}(\cdot,t),x_{2}(\cdot,t),\partial_{t}x_{2}(\cdot,t))^{\intercal}$ and $\mathcal{X}=\{\mathbb{R}^{2}\times L^{2}([-\alpha,0])\times H^{1}([-\alpha,0])\times L^{2}([0,\beta])\times H^{1}([0,\beta])\;|\;\xi(t)=x_{1}(0,t)=x_{2}(0,t)\}$. $\bm{x}(\cdot,t)\in\mathcal{X}$ as initially discussed in [7]. By using the deflection and gradient at $z=0$ as Cauchy data, D’Alemberts’s formula yields the solutions
$x_{i}(z,t)\!=\!\frac{1}{2}\!\left(x_{i}(0,t-z)\!+\!x_{i}(0,t+z)\right)\!+\!\frac{1}{2}\!\int\limits_{t-z}^{t+z}\partial_{z}x_{i}(0,\tau)\,\mathrm{d}\tau$ (6)
for $i=1,2$ (c.f. [7, 4]), which reads
$x_{i}(z,t)=C_{z}x_{i}(0,t)+\frac{1}{s}S_{z}\partial_{z}x_{i}(0,t)$ (7)
in the algebraic setting, where $C_{z}\coloneqq\cosh(\partial_{t}z)=\frac{1}{2}(\sigma^{z}+\sigma^{-z})$ and $S_{z}\coloneqq\sinh(\partial_{t}z)=\frac{1}{2}(\sigma^{z}-\sigma^{-z})$.
II-D Parametrisation
Starting at the interior mass, by using the tuple
$\begin{pmatrix}{y}_{1}(t),{y}_{2}(t)\end{pmatrix}^{\intercal}=\begin{pmatrix}\xi(t),\partial_{z}x_{1}(0,t)+\partial_{z}x_{2}(0,t)\end{pmatrix}^{\intercal}$ (8)
with $y_{1}\in H^{2}(\mathbb{R})$ and $y_{2}\in L^{2}(\mathbb{R})$ we directly obtain $\xi(t)=x_{1}(0,t)=x_{2}(0,t)={y}_{1}(t)$ as well as
$\partial_{z}x_{1}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)-m\partial_{t}^{2}{y}_{1}(t)\right)$ (9a)
$\partial_{z}x_{2}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)+m\partial_{t}^{2}{y}_{1}(t)\right)$ (9b)
from (5b). With the Cauchy data known, (7) then yields the profiles, from which the input parameterisation
$\begin{pmatrix}u_{1}(t)\\ u_{2}(t)\\ \end{pmatrix}=\begin{pmatrix}\frac{m}{2}\partial_{t}^{2}C_{\alpha}+\partial_{t}S_{\alpha}&-\frac{1}{2}C_{\alpha}\\ \frac{m}{2}\partial_{t}^{2}C_{\beta}+\partial_{t}S_{\beta}&\frac{1}{2}C_{\beta}\end{pmatrix}\begin{pmatrix}y_{1}(t)\\ y_{2}(t)\\ \end{pmatrix}$ (10)
is obtained via (5c) and (5d) showing that (8) is a flat output of the system (4)-(5).
II-E Stating a controller form
As known from, e.g [4], the problem to be solved for stating a controller form is to find a particular realisation of the input-output relation (10). A possible way to do so is to solve (10) for the highest derivatives of the largest predictions of $y_{1}(t)$ as well as $y_{2}(t)$ and to check for which initial conditions the resulting DDE system is well posed.
However, taking the coefficients of the highest derivatives of the largest predictions in each column on the right hand-side of (10), the resulting leading column coefficient matrix
$\operatorname{lccm}_{\rho,s}{\mathbf{H}}=\frac{1}{4}\begin{pmatrix}0&0\\ m&1\\ \end{pmatrix}$ (11)
does not have full rank. This implies that (10) cannot directly be written as an explicit DDE system thus negating our approach. Therefore, more steps have to be taken for MIMO systems, which will be discussed in Section IV after we give a definition of the desired controller form for this case.
III A hyperbolic controller form
In this section we define a general controller form for hyperbolic ODE-PDE MIMO systems and state several specific variants of the latter which may be helpful in different settings.
III-A Definition
By directly applying the extension of the SISO controller form by a transport system from (3) to the MIMO controller form (2), for $i\in\{1,\dotsc,m\}$, we obtain the subsystems
$\dot{\eta}^{\mathrm{O}}_{i,1}(t)$ $=\eta^{\mathrm{O}}_{i,2}(t)$ (12a)
$\;\;\vdots$
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}-1}(t)$ $=\eta^{\mathrm{O}}_{i,\nu_{i}}(t)$ (12b)
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=\eta^{\mathrm{P}}_{i}(\check{\tau}_{i},t)$ (12c)
$\partial_{t}\eta^{\mathrm{P}}_{i}(\tau,t)$ $=\partial_{\tau}\eta^{\mathrm{P}}_{i}(\tau,t),\quad\tau\in[\check{\tau}_{i},\hat{\tau}_{i})$ (12d)
$\begin{split}\eta^{\mathrm{P}}_{i}(\hat{\tau}_{i},t)&=-\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{a^{\star}_{i,j}}\eta^{\mathrm{P}}_{j}(\cdot,t)+\bm{a}_{i,j}^{\intercal}\eta^{\mathrm{O}}_{j}(t)\\ &\qquad\qquad+\operatorname{\mathcal{S}}_{b_{i,j}^{\star}}\!*\,u_{j}(t),\end{split}$ (12e)
with state $\bm{\eta}(\cdot,t)\in\mathbb{R}^{\nu_{\Sigma}}\times\prod_{i=1}^{m}L^{2}([\check{\tau}_{i},\hat{\tau}_{i}])$. Herein, the convolutions $\operatorname{\mathcal{S}}_{b_{i,j}^{\star}}\!*\,u_{j}(t)=\int_{\underline{\tau\!}}^{\bar{\tau}}u_{j}(t-\tau)\,\mathrm{d}b_{i,j}^{\star}(\tau)$ with the $b_{i,j}^{\star}$ to be further specified form the input map $\mathbf{B}^{\star}=[\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}]$ which is invertible w.r.t. convolution.
In (12), each block consist of a transport system whose output feeds into an integrator chain. However, the input of each block’s transport system now depends on the system input as well as the PDE and ODE states of all blocks. In the following, we will derive several special cases of the form (12) by specifying the functions $b_{i,j}^{\star}$ that define $\mathbf{B}^{\star}$.
III-B A classic controller form
To recover a classic form without any input derivatives, predictions or delays, we restrain the functions to $b_{i,j}^{\star}(t)=b_{i,j}H(t)$ with the Heavyside function $H(t)$, resulting in
$\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}\!*\,{u_{j}}(t)=\bm{b}_{i}^{\intercal}\bm{u}(t)$ (13)
for $i\in\{1,\dotsc,{m}\}$ after evaluating the convolutions.
III-C Generalised controller forms
III-C1 With input derivatives
To allow for input derivatives to occur in (12e), we set $b_{i,j}^{\star}(t)=\sum_{k=0}^{D_{i,j}}b_{k}H^{(k)}(t)$, yielding
$\sum\limits_{j=1}^{m}\operatorname{\mathcal{S}}_{b^{\star}_{i,j}}\!*\,{u_{j}}(t)=\sum\limits_{j=1}^{m}\sum_{k=0}^{D_{i,j}}b_{i,j,k}u_{j}^{(k)}(t)$ (14)
where the $D_{i,j}$ denote the highest derivate order of $u_{j}(t)$ in the equations of subsystem $i$. Owing to the derivatives of the $u_{j}\in L^{2}(\mathbb{R}^{+})$, we propose the term discontinuous controller form for the subtype (14).
III-C2 With input predictions
For this case, we have $b_{i,j}^{\star}\in BV([-\bar{\tau}_{i,j},0])$, where the $\bar{\tau}_{i,j}$ denote the biggest predictions with which the input $u_{j}(t)$ appears in the subsystem $i$. As such a system depends on future inputs, we propose the term non-causal controller form for this case.
III-C3 With input predictions and delays
For a more general case, we allow $b_{i,j}^{\star}\in BV([-\bar{\tau}_{i,j},-\underline{\tau\!}_{i,j}])$, where the $\underline{\tau\!}_{i,j}$ denote the biggest delays of input $u_{j}(t)$. As this form only yields a sensible state if the system is defined for $t\in\mathbb{R}$ instead of $\mathbb{R}^{+}$, we propose the term quasi controller form.
IV Computation of the controller form
This section contains the main result of the contribution, in which we present an algorithm to compute the controller form for a specific class of systems.
IV-A System Class
We discuss the linear first order hyperbolic MIMO system [^3] Note that the system from Section II-C can easily be transformed into this more general form which also allows for unsymmetrical predictions and delays in contrast to the more compact second order formulation.
$\dot{\bm{\xi}}(t)$ $=\bm{F}\bm{\xi}(t)+\bm{B}\bm{x}^{-}(0,t)$ (15a)
$\partial_{t}\bm{x}(z,t)$ $=\bm{\Lambda}\partial_{z}\bm{x}(z,t),\quad z\in(0,1)$ (15b)
$\bm{x}^{+}(0,t)$ $=\bm{Q}_{0}\bm{x}^{-}(0,t)+\bm{C}\bm{\xi}(t)$ (15c)
$\bm{x}^{-}(1,t)$ $=\bm{Q}_{1}\bm{x}^{+}(1,t)+\bm{u}(t)$ (15d)
as depicted in Figure 1, where $\bm{\xi}(t)\in\mathbb{R}^{n}$, $\bm{x}^{-}(\cdot,t)\in L^{2}([0,1])^{{n^{-}}}$ and $\bm{x}^{+}(\cdot,t)\in L^{2}([0,1])^{n^{+}}$ denote the boundary system’s as well as the backward and forward transport systems’ states. This model is also known as a PDE-ODE system in the backstepping literature, cf. i.e. [8].
For (15), we assume decoupled PDEs with $\bm{\Lambda}=\mathrm{diag}(\mathbf{\Lambda}^{-},-\mathbf{\Lambda}^{+})$ where $\mathbf{\Lambda}^{\pm}=\mathrm{diag}(\lambda_{1}^{\pm},\dotsc,\lambda_{n^{\pm}}^{\pm})$, $\lambda_{i}^{\pm}>\lambda_{i+1}^{\pm}>0$, the pair $(\bm{F},\bm{B})$ controllable and $\operatorname{rank}\bm{Q}_{0}={n^{+}}$ as well as $\operatorname{rank}\bm{Q}_{1}={n^{-}}$.
Figure 1: Structure of the assumed system class
IV-B Parametrisation
As (15a) is controllable, a flatness-based parametrisation in the form of
$\bm{\xi}(t)$ $=\mathbf{N}(\partial_{t})\bm{y}(t)$ (16a)
$\bm{x}^{-}(0,t)$ $=\mathbf{D}(\partial_{t})\bm{y}(t)$ (16b)
can always be given. Furthermore, if the components of the flat output $\bm{y}(t)$ are chosen as the outputs of each block’s integrator chain of the systems’ MIMO controller form (2), the parametrisation matrix $\mathbf{D}(\partial_{t})$ is column reduced and $\operatorname{coldeg}_{i}\mathbf{N}(\partial_{t})<\operatorname{coldeg}_{i}\mathbf{D}(\partial_{t}),\;i\in\{1,\dotsc,m\}$ holds for the column degrees.
Regarding the PDE subsystem, with $\bm{\xi}(t)$ as well as $\bm{x}^{-}(0,t)$ known, (15c) yields $\bm{x}^{+}(0,t)=\left(\mathbf{Q}_{0}\mathbf{D}(\partial_{t})+\mathbf{C}\mathbf{N}(\partial_{t})\right)\bm{y}(t)$ and thus the boundary values required to solve (15b) up to $z=1$. This yields
$\bm{x}^{-}(1,t)$ $=\mathbf{P}(\sigma,t)\bm{x}^{-}(0,t)$ (17)
$\bm{x}^{+}(1,t)$ $=\mathbf{\Delta}(\sigma,\partial_{t})\bm{x}^{+}(0,t)$ (18)
with the prediction matrix $\mathbf{P}(\sigma)=\operatorname{diag}(\sigma^{\tau_{1}^{-}},\dotsc,\sigma^{\tau_{{n^{-}}}^{-}})$ and the delay matrix $\mathbf{\Delta}(\sigma)=\operatorname{diag}(\sigma^{-\tau_{1}^{+}},\dotsc,\sigma^{-\tau_{{n^{+}}}^{+}})$ where $\tau_{i}^{\pm}=\frac{1}{\lambda_{i}^{\pm}}$. Finally, combining Equations 16a, 16b, 17 and 18 with (15d) yields the matrix
$\mathbf{H}(\sigma,\partial_{t})\!=\!\mathbf{P}(\sigma)\mathbf{D}(\partial_{t})-\mathbf{Q}_{1}\mathbf{\Delta}(\sigma)\left(\mathbf{Q}_{0}\mathbf{D}(\partial_{t})+\mathbf{C}\mathbf{N}(\partial_{t})\right)$ (19)
for the parametrisation
$\bm{u}(t)=\mathbf{H}(\sigma,\partial_{t})\bm{y}(t)$ (20)
of the system input by the flat output of the ODE.
As was already to be expected from the motivational example’s parametrisation (10), (19) now structurally shows that $\mathbf{\mathbf{P}}(\sigma)$ and $\mathbf{\Delta}(\sigma)$ introduce all predictions and delays that occur in the system into every column of $\mathbf{H}(\sigma,\partial_{t})$ which poses a problem. Next, we will extend the notion of polynomials in $\partial_{t}$ to incorporate shifts and remedy this inconvenience.
IV-C Generalised polynomials
With the set $T=\{\tau_{1}^{-},\dotsc,\tau_{n^{-}}^{-},\tau_{1}^{+},\dotsc,\tau_{n^{+}}^{+},\}$ of all unique predictions and shifts in the system, by interpreting the members of $\Sigma=\{\sigma^{\tau},\sigma^{-\tau}|\tau\in T\}$ and $\partial_{t}$ as independent generators, we conceive the entries $h_{i,j}(\sigma,\partial_{t})$ of $\mathbf{H}(\sigma,\partial_{t})$ to be from $\mathcal{R}_{\mathrm{p}}=\mathbb{R}[\Sigma,\partial_{t}]$, the space of generalised polynomials in $\sigma$ with real exponents, whose coefficients are polynomials in $\partial_{t}$ with coefficients from $\mathbb{R}$. Furthermore, for the elements
$r(\sigma,\partial_{t})=\sum_{i=1}^{n}a_{i}(\partial_{t})\sigma^{\alpha_{i}}\in\mathcal{R}_{\mathrm{p}}$ (21)
with $a_{i}(\partial_{t})\in\mathbb{R}[\partial_{t}]$, the ordering $\alpha_{i}>\alpha_{i+1}$, for $i\in\{1,\dotsc,n-1\}$ is assumed. To ease working on the polynomials (21), we define the pseudo degrees
$\deg_{\sigma}^{+}r(\sigma,\partial_{t})\coloneq\max_{i\in\{1,\dotsc,n\}}\alpha_{i},\;\deg_{\sigma}^{-}r(\sigma,\partial_{t})\coloneq\min_{i\in\{1,\dotsc,n\}}\alpha_{i}$
as well as the degree [^4] This choice is motivated by the case of Laurent polynomials, which posses positive and negative integer exponents and form an euclidean ring for the given degree definition, cf. [9, 10].
$\deg_{\sigma}r(\sigma,\partial_{t})\coloneq\deg_{\sigma}^{+}r(\sigma,\partial_{t})-\deg_{\sigma}^{-}r(\sigma,\partial_{t}),$ (22)
yielding the upper and lower boundary as well as the diameter of the interval, respectively. Next, we discuss how these intervals can be reduced in a systematic way.
IV-D Shift interval reduction
IV-D1 Preliminary Transformations
Lemma 1
By permuting the input and/or flat output, $\mathbf{H}(\sigma,\partial_{t})$ can be sorted, such that the degrees w.r.t. $\sigma$ in each column are ascending and the highest degrees w.r.t. $\partial_{t}$ appear on the main diagonal, thus with $j\in\{1,\dotsc,m\}$
$\deg_{\sigma}{h}_{i,j}(\sigma,\partial_{t})$ $\leq\deg_{\sigma}{h}_{i+1,j}(\sigma,\partial_{t}),$ (23a)
$\deg_{\partial_{t}}{h}_{l,j\neq}(\sigma,\partial_{t})$ $\leq\deg_{\partial_{t}}{h}_{l,l}(\sigma,\partial_{t})$ (23b)
hold for $i\in\{1,\dotsc,m-1\}$ and $l\in\{1,\dotsc,m\}$.
Proof:
As from (19), there are equal predictions and delays in each row of $\mathbf{H}(\sigma,\partial_{t})$ due to the diagonal structure of $\mathbf{P}(\sigma)$ and $\mathbf{\Delta}(\sigma)$. Then, by (22) the same follows for their corresponding degrees which is why the ordering (23a) is easily obtained by permuting the flat output.
Now, from the parametrisation (16) we have that $\operatorname{coldeg}\mathbf{N}(\partial_{t})<\operatorname{coldeg}\mathbf{D}(\partial_{t})$, which is why $\mathbf{C}$ has no effect regarding the degrees of $\partial_{t}$} in (19) and also that $\mathbf{D}(\partial_{t})$ is column reduced. Since this property is preserved under row transformations over $\mathbb{R}$ and $\mathbf{Q}_{0},\mathbf{Q}_{1}$ have full rank, we obtain that $\mathbf{H}(\sigma,\partial_{t})$ is column reduced w.r.t. $\partial_{t}$. Finally, as the elements in each row have equal degree w.r.t. $\sigma$, (23b) is achieved by permuting the input. ∎
IV-D2 Goal
As discussed in the motivation, to obtain the controller form (12), the input parametrisation (20) must be solved for the highest derivatives of the flat output. Furthermore, to introduce a HCF state in a direct fashion, these must occur together with the largest predictions as well as largest delays. In the algebraic setting, this corresponds to the composition of taking the LCCMs w.r.t. $\sigma$ and $\partial_{t}$ of $\mathbf{H}(\sigma,\partial_{t})$ to commute:
$\operatorname{lccm}_{s}(\operatorname{lccm}^{+}_{\sigma}\mathbf{H}(\sigma,\partial_{t}))$ $=\operatorname{lccm}^{+}_{\sigma}(\operatorname{lccm}_{s}\mathbf{H}(\sigma,\partial_{t}))$ (24a)
$\operatorname{lccm}_{s}(\operatorname{lccm}_{\sigma}\mathbf{H}(\sigma,\partial_{t}))$ $=\operatorname{lccm}_{\sigma}(\operatorname{lccm}_{s}\mathbf{H}(\sigma,\partial_{t}))$ (24b)
and the result having full rank. The condition for $\deg_{\sigma}^{-}$ then follows from (22). These requirements are trivially satisfied if the degree conditions
$\deg^{+}_{\sigma}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg^{+}_{\sigma}{h}_{i,i}(\sigma,\partial_{t})$ (25a)
$\deg_{\sigma}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg_{\sigma}{h}_{i,i}(\sigma,\partial_{t})$ (25b)
$\deg_{\partial_{t}}{h}_{i,j\neq i}(\sigma,\partial_{t})$ $\leq\deg_{\partial_{t}}{h}_{i,i}(\sigma,\partial_{t})$ (25c)
hold for all $i,j\in\{1,\dotsc,{m}\}$, since all LCCMs in this case are diagonal. Furthermore, as by Lemma 1, (25a),(25b) for $j\leq i$ as well as (25c) for $j\in\{1,\dotsc,m\}$ are satisfied, the elements $h_{i,i}(\sigma,\partial_{t})$ of each column are now used to reduce the (pseudo) degrees of the entries below them, such that all
$\bar{h}_{j,i}(\sigma,\partial_{t})=h_{i,j}(\sigma,\partial_{t})-q_{j,i}(\sigma,\partial_{t})h_{i,i}(\sigma,\partial_{t})$ (26)
for $i,j\in\{1,\dotsc,m\},j>i$ satisfy (25). Next, we discuss a generalised polynomial long division (GPLD) algorithm to perform such a reduction for a single element.
IV-D3 Generalised polynomial long division
For now, we treat all entries in (26) as elements of $\mathcal{R}_{\mathrm{r}}=\mathbb{R}(\partial_{t})\left[\Sigma\right]$ which allows for rational expressions in $\partial_{t}$ as coefficients and therefore enables us to develop a polynomial long division scheme regarding $\sigma$.
Lemma 2
Let $x(\sigma,\partial_{t}),y(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}},\;y(\sigma,\partial_{t})\neq 0$ be generalised polynomials with rational coefficients where $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$. Then there exist $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}}$ such that
$x(\sigma,\partial_{t})=q(\sigma,\partial_{t})y(\sigma,\partial_{t})+r(\sigma,s)$ (27)
with $\deg_{\sigma}r(\sigma,\partial_{t})<\deg_{\sigma}y(\sigma,\partial_{t})$ as well as $\deg^{+}_{\sigma}r(\sigma,\partial_{t})<\deg^{+}_{\sigma}y(\sigma,\partial_{t})$ and $\deg^{-}_{\sigma}r(\sigma,\partial_{t})>\deg^{-}_{\sigma}y(\sigma,\partial_{t})$.
Proof:
As the following steps are valid for arbitrary coefficients, the explicit mention of $\partial_{t}$ will be omitted. In the general case we have $\deg_{\sigma}x(\sigma)>\deg_{\sigma}y(\sigma)$ as well as $\deg^{+}_{\sigma}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)$, $\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma)$ or both. Since the exponents of the polynomials $x(\sigma,\partial_{t})=\sum_{i=1}^{n}a_{i}(\partial_{t})\sigma^{\alpha_{i}}$, $y(\sigma,\partial_{t})=\sum_{i=1}^{m}b_{i}(\partial_{t})\sigma^{\beta_{i}}$ are ordered, we use
$q(\sigma)=\begin{cases}\frac{a_{1}\sigma^{\alpha_{1}}}{b_{1}\sigma^{\beta_{1}}},&\text{ if }\deg_{\sigma}^{+}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)\\ \frac{a_{n}\sigma^{\alpha_{n}}}{b_{m}\sigma^{\beta_{m}}},&\text{ if }\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma),\end{cases}$ (28)
where in the case that both are true, the upper branch is taken. Assuming the upper branch was taken, this yields
$\begin{split}r(\sigma)=x(\sigma)-q(\sigma)y(\sigma)=\sum\limits_{i=2}^{n}a_{i}\sigma^{\alpha_{i}}-\frac{a_{1}\sigma^{\alpha_{1}}}{b_{1}\sigma^{\beta_{1}}}\sum\limits_{i=2}^{m}b_{i}\sigma^{\beta_{i}},\end{split}$
for which we find that
$\deg^{+}_{\sigma}r(\sigma)=\max\left\{\alpha_{2},\alpha_{1}-(\beta_{1}-\beta_{2})\right\}<\deg^{+}_{\sigma}x(\sigma)$
as well as $\deg^{-}_{\sigma}r(\sigma)=\min\{\alpha_{n},\alpha_{1}-(\beta_{1}-\beta_{m})\}$ in which we identify $\deg_{\sigma}y(\sigma)=\beta_{1}-\beta_{m}$ and use $\alpha_{1}=\deg_{\sigma}x(\sigma)+\alpha_{n}$ as well as $\kappa=\deg_{\sigma}x(\sigma)-\deg_{\sigma}y(\sigma)\geq 0$ to obtain
$\begin{split}\deg^{-}_{\sigma}r(\sigma)=\min\{&\alpha_{n},\alpha_{n}+\kappa\}=\alpha_{n}=\deg^{-}_{\sigma}x(\sigma).\end{split}$
This shows that the predictions in $x(\sigma)$ can be reduced without affecting the delays. Regarding the overall degree, we have $\deg_{\sigma}r(\sigma)=\max\{\alpha_{2}-\alpha_{n},\alpha_{1}-\alpha_{n}-(\beta_{1}-\beta_{2})\}$ which yields $\deg_{\sigma}r(\sigma)<\deg_{\sigma}x(\sigma)$. Hence, $x(\sigma)$ can be reduced in a manner such that the iteration
$r_{k}(\sigma)=r_{k-1}(\sigma)-q_{k}(\sigma)y(\sigma)$
with $r_{0}(\sigma)=x(\sigma)$ and $q_{k}(\sigma)=\frac{c_{k,1}\sigma^{\gamma_{k,1}}}{b_{1}\sigma^{\beta_{1}}}$ by $r_{k}(\sigma)=\sum_{i=1}^{n}c_{k,i}\sigma^{\gamma_{k,i}}$ will for some $k^{+}$ terminate with $\deg^{+}_{\sigma}r_{k^{+}}(\sigma)<\deg^{+}_{\sigma}y(\sigma)$ since $x(\sigma)$ only has a finite amount of summands.
Now, (if required) the second branch of (28) is taken and similar computations yield that restarting the iteration with $r_{0}(\sigma)=r_{k^{+}}(\sigma)$ and choosing $q_{k}(\sigma)=\frac{c_{k,1}\sigma^{\gamma_{k,1}}}{b_{m}\sigma^{\beta_{m}}}$ will yield $\deg^{-}_{\sigma}r_{k^{-}}(\sigma)>\deg^{-}_{\sigma}y(\sigma)$ after $k^{-}$ more steps. Summing up all $q_{k}(\sigma)$ from both iterations then yields the resulting “quotient” $q(\sigma)$ with common denominator $d=b_{1}^{k^{+}}b_{m}^{k^{-}}$. ∎
With Lemma 2, the shift degrees are taken care of. However, the coefficients of $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})$ are rational expressions in $\partial_{t}$. Furthermore, in the analytic setting, their denominators can be interpreted as functions that their numerator polynomials are applied to. This results in piecewise continuous functions which do not necessarily have compact support. In other words: They generally do not correspond to differential or (distributed) delay operators with finite shift amplitude from $\mathcal{E}^{\prime}$. However, thanks to the Palay-Wiener theorem, if their Laplace transforms are entire functions, they indeed are such operators, cf. [11, Thm. 7.3.1].
Lemma 3
Let $x(\sigma,\partial_{t}),y(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{p}},\;y(\sigma,\partial_{t})\neq 0$ be generalised polynomials with polynomial coefficients where $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$, as well as $\deg_{\partial_{t}}x(\sigma)\leq\deg_{\partial_{t}}y(\sigma)$. Then there exist $q^{\star}(\sigma,\partial_{t}),r^{\star}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{w}}=\mathcal{R}_{\mathrm{r}}\cap\mathcal{E}^{\prime}$ such that
$x(\sigma,\partial_{t})=q^{\star}(\sigma,\partial_{t})y(\sigma,\partial_{t})+r^{\star}(\sigma,s)$ (29)
with $\deg_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg_{\sigma}y(\sigma,\partial_{t})$, $\deg^{+}_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg^{+}_{\sigma}y(\sigma,\partial_{t})$ and $\deg^{-}_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg^{-}_{\sigma}y(\sigma,\partial_{t})$.
Proof:
This proof is heavily inspired by [12, 9] and only adapted to the current setup. First, we obtain $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}}$ by means of Lemma 2 and extract its normalised common denominator $d(\partial_{t})\in\mathbb{R}[\partial_{t}]$ such that $q(\sigma,\partial_{t})=\frac{\tilde{q}(\sigma,\partial_{t})}{d(\partial_{t})}$ with $\tilde{q}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{p}}$. Next, the factorisation $d(\partial_{t})=\prod_{k=1}^{D}(\partial_{t}-s_{k})$ with the roots $s_{k}$ and $D=\deg d(\partial_{t})$ are used together with $\hat{q}_{0}(s)=\mathcal{L}\{\tilde{q}(\sigma,\partial_{t})\}$ in the iteration
$\hat{q}_{k}(s)=\frac{\hat{q}_{k-1}(s)-\hat{q}_{k-1}(s_{k})}{s-s_{k}},$ (30)
which introduces a nominator zero for every pole from $q(\sigma,\partial_{t})$ it adds. Thus, $q_{D}(s)$ obtained after $D$ iterations is an entire function on $\mathbb{C}$ and therefore $q^{\star}(\sigma,\partial_{t})=\mathcal{L}^{-1}\{q_{D}(s)\}\in\mathcal{R}_{\mathrm{w}}$ by [11, Thm. 7.3.1].
Now, due to the structure of (30) we can justify the separation $q^{\star}(\sigma,\partial_{t})=q(\sigma,\partial_{t})+p(\partial_{t})$ with $p(\partial_{t})\in\mathbb{R}(\partial_{t})$. Therefore, we obtain the degree conditions from Lemma 2 with equal signs for the new remainder $r^{\star}(\sigma,\partial_{t})$. ∎
IV-D4 Algorithm
Equipped with a suitable way to reduce the degrees of its elements by means of the GPLD from Lemma 3, we now reduce the degrees in the entire matrix $\mathbf{H}(\sigma,\partial_{t})$ by applying a series of row transformations.
Algorithm 1 Shift reduction in $\mathbf{H}(\sigma,\partial_{t})$
alg]alg:shift_red
1: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]function compute_transform($\mathbf{X},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
2:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow\mathbf{I}_{{n^{-}}}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
3:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $i\in\{j+1,\dotsc,{n^{-}}\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
4:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] // From Lemma 3: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
5:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $q^{\star}_{i,j}\leftarrow$ GPLD($x=\mathbf{X}[i,j],y=\mathbf{X}[j,j]$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
6:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}[i,j]\leftarrow-q^{\star}_{i,j}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
7:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] return $\mathbf{L}_{c}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
8:
9: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]procedure reduce_shifts($\mathbf{H}$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
10:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{H}\leftarrow\mathbf{H}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
11:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $i\in\{2,\dotsc,{n^{-}}\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
12:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $j\in\{1,\dotsc,{n^{-}}-1\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
13:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow$ Compute transform($\bar{}\mathbf{H},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
14:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{H}\leftarrow\mathbf{L}_{c}\mathbf{H}$, $\mathbf{L}\leftarrow\mathbf{L}_{c}\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
15:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] return $\bar{}\mathbf{H}$$\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
Lemma 4
The matrix $\bar{}\mathbf{H}(\sigma,\partial_{t})=\mathbf{L}(\sigma,\partial_{t})\mathbf{H}(\sigma,\partial_{t})$ obtained by LABEL:alg:shift_red satisfies the requirements (24).
(25).
Proof:
Starting in the second row, the degrees in each column are corrected from left to right, where each of these correction steps will introduce new undesired shifts to its left in the rows below its diagonal element. Therefore, the procedure then continues to the next row, until no more new shifts are introduced as the last row is reached. ∎
IV-E Hyperbolic MIMO controller form
Applying LABEL:alg:shift_red to $\mathbf{H}(\sigma,\partial_{t})$, we obtain
$\mathbf{L}^{-1}(\sigma,\partial_{t})\bm{u}(\partial_{t})=\bar{}\mathbf{H}(\sigma,\partial_{t})\bm{y}(\partial_{t})$ (31)
with $\hat{\mathbf{H}}=\operatorname{lccm}_{s,\sigma}\bar{}\mathbf{H}(\sigma,\partial_{t})$ having full rank. Thus, by means of the separation $\bar{}\mathbf{H}(\sigma,\partial_{t})=\hat{}\mathbf{H}\mathbf{K}(\sigma,\partial_{t})+\tilde{}\mathbf{H}(\sigma,\partial_{t})$ where $\mathbf{K}(\sigma,\partial_{t})=\operatorname{diag}(\partial_{t}^{\nu_{1}}\sigma^{\hat{\tau}_{1}},\dotsc,\partial_{t}^{\nu_{m}}\sigma^{\hat{\tau}_{m}})$ with the highest derivative and prediction orders $\nu_{i}=\deg_{\partial_{t}}\bar{h}_{i,i}(\sigma,\partial_{t})$ and $\hat{\tau}_{i}=\deg^{+}_{\sigma}\bar{h}_{i,i}(\sigma,\partial_{t})$ we obtain the explicit DDE system
$\mathbf{K}(\sigma,\partial_{t})\bm{y}(\partial_{t})=\hat{\mathbf{H}}^{-1}\!\left(\mathbf{L}^{-1}(\sigma,\partial_{t})\bm{u}(\partial_{t})\!-\tilde{\mathbf{H}}(\sigma,\partial_{t})\bm{y}(\partial_{t})\right)$ (32)
which reads
$y_{i}^{(\nu_{i})}(t+\hat{\tau}_{i})=-\mathcal{A}_{i}(t)*\bm{y}(t)+\mathcal{B}_{i}(t)*\bm{u}(t)$ (33)
for $i=\{1,\dotsc,m\}$ in the analytic setting. Herein, the elements of $\mathcal{A}_{i},\mathcal{B}_{i}$ by construction have compact support, which is why the convolutions will yield definite integrals whose intervals depend on the degrees of $\tilde{\mathbf{H}}(\sigma,\partial_{t})$ and $\mathbf{L}^{-1}(\sigma,\partial_{t})$ w.r.t $\sigma$. Thus, (33) states a well posed initial value problem if the components $y_{i}(t+\tau_{i})$ of the flat output are known for $\tau_{i}\in[\check{\tau}_{i},\hat{\tau}_{i}]$ with $\check{\tau}_{i}=\deg^{-}_{\sigma}\bar{h}_{i,i}(\sigma,\partial_{t})$ and $\bm{u}$ is known on $\mathbb{R}$.
Hence, we introduce the HCF state
$\eta^{\mathrm{O}}_{i,j}(t)$ $\coloneq y_{i}^{(j-1)}(t-\check{\tau}_{i}),\;j\in\{1,\dotsc,\nu_{i}\}$ (34a)
$\eta^{\mathrm{P}}_{i}(\tau,t)$ $\coloneq y_{i}^{(\nu_{i})}(t+\tau),\;\tau\in[\check{\tau}_{i},\hat{\tau}_{i}]$ (34b)
for $i\in\{1,\dotsc,m\}$ which is equivalent to the flat output on these intervals. Substituting [^5] Some expressions in $\bm{y}(t)$ may remain as e.g. the values $y_{i}^{(d)}(t+\tau_{i})$ are not part of the HCF state $\bm{\eta}(\cdot,t)$ for $p<\nu_{i}$, cf. (12). However, these expressions can be easily eliminated by means shown e.g. in [6, Lemma 5]. the definitions (34) in (33) then allows to identify the functions $a^{\star}_{i,j}(t),b^{\star}_{i,j}(t)$ as well as weights $\bm{a}_{i}$ in (12), yielding the HCF of system (15).
Lemma 5
The type of HCF obtained is a quasi controller form, (cf. Section III-C3), with only input predictions and delays but no input derivatives.
Proof:
The proof of Lemma 2 yields $\deg_{\partial_{t}}q(\sigma,\partial_{t})<0$ due to the structure of $d(\partial_{t})$. Then by Lemma 3, $q^{\star}(\sigma,\partial_{t})$ has non-positive degree as do all entries in $\mathbf{L}_{i}(\sigma,\partial_{t})$ from LABEL:alg:shift_red. Now, as the inverse of the triangular $\mathbf{L}(\sigma,\partial_{t})$ is computed by forward substitution, all $\bar{q}_{i,j}(\sigma,\partial_{t})$ therein are linear combinations of the $q_{i,j}^{\star}(\sigma,\partial_{t})$, hence, $\deg_{\partial_{t}}\bar{q}_{i,j}(\sigma,\partial_{t})\leq 0$. Therefore, the convolution with $\mathcal{B}_{i}$ yields no derivatives. ∎
V Reduction of the benchmark system
Consider again the system from Section II-C with $\alpha=\pi$, $\beta=10$ and $m=1$ substituted into (10) yielding
$\mathbf{H}(\sigma,\partial_{t})=\frac{1}{4}\begin{pmatrix}a_{1}(\partial_{t})\sigma^{\pi}+a_{2}(\partial_{t})\sigma^{-\pi}&-\sigma^{\pi}-\sigma^{-\pi}\\ b_{1}(\partial_{t})\sigma^{10}+b_{2}(\partial_{t})\sigma^{-10}&\sigma^{10}+\sigma^{-10}\\ \end{pmatrix}$
with $a_{1}(\partial_{t})=b_{1}(\partial_{t})=\partial_{t}^{2}+2\partial_{t}$ and $a_{2}(\partial_{t})=b_{2}(\partial_{t})=\partial_{t}^{2}-2\partial_{t}$ for which $\operatorname{lccm}_{\sigma,\partial_{t}}\mathbf{H}(\sigma,\partial_{t})=\operatorname{lccm}_{\partial_{t},\sigma}\mathbf{H}(\sigma,\partial_{t})$ is of rank $1$ where $2$ would be needed.
Now, regarding Section IV-D1, the rows are already correctly ordered since $\pi<10$ and also the columns do not need to be reordered since the coefficients in each column are of the same degree w.r.t. $\partial_{t}$. Continuing with Section IV-D2, the conditions (25) fail for $h_{2,1}(\sigma,\partial_{t})$ as e.g. $\deg_{\sigma}h_{2,1}(\sigma,\partial_{t})=20>\deg_{\sigma}h_{1,1}(\sigma,\partial_{t})=2\pi$. Therefore, by means of Lemma 2 the quotient
$\begin{aligned} q_{1,2}&(\sigma,\partial_{t})\!=\!\sigma^{10-\pi}\!+\!\frac{2-\partial_{t}}{\partial_{t}+2}\sigma^{10-3\pi}\!-\!\frac{\partial_{t}+2}{\partial_{t}-2}\sigma^{-10+3\pi}\!+\!\sigma^{-10+\pi}\end{aligned}$
is obtained, where one iteration was needed for each case in (28). Next, Lemma 3 yields the correction term $p(\partial_{t})=\frac{16e^{-20+6\pi}}{\left(\partial_{t}^{2}-4\right)}$ and thus $q_{2,1}^{\star}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{w}}$. Using the latter in (26) then yields the reduced element
$\begin{aligned} \bar{h}_{1,2}(\sigma,\partial_{t})=\!-\frac{16\partial_{t}}{\left(\partial_{t}-2\right)e^{20-6\pi}}\sigma^{\pi}\!+\frac{\partial_{t}\left(\partial_{t}^{2}+4\partial_{t}+4\right)}{\partial_{t}-2}\sigma^{-10+4\pi}\\ \!+\frac{\partial_{t}\left(\partial_{t}^{2}-4\partial_{t}+4\right)}{\partial_{t}+2}\sigma^{10-4\pi}\!-\frac{16\partial_{t}}{\left(\partial_{t}+2\right)e^{20-6\pi}}\sigma^{-\pi}.\end{aligned}$
As the requirements for the second column are already satisfied, this concludes the reduction and we obtain $\hat{}\mathbf{H}=\operatorname{lccm}_{\partial_{t},\sigma}\bar{}\mathbf{H}(\sigma,\partial_{t})=\operatorname{diag}(1,-2)$.
Python code showing the complete computation of the HCF will be made available at a later time.
VI Concluding remarks
This contribution establishes a HCF for MIMO systems with dynamic boundaries and proposes a flatness based algorithm to compute said form for a specific class of systems. While the algebraic approach taken to obtain this result is limited to linear systems, it enables a systematic procedure which lends well to an implementation using computer algebra. However, the insight gained from an analytic interpretation of the algebraic procedure allows for the technique to be extended to nonlinear boundary systems, which will be discussed in a forthcoming publication. Furthermore, the obtained HCF contains input delays, which will result in a dynamic extension, thus further analysis is required to obtain a form without delays as this would enable the application of quasi-static state feedback. Finally, the system class (15) is lacking in-domain couplings, as these were recently addressed in [13] for the MIMO case without boundary dynamics, the next logical step is to combine the two approaches.